• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Iodice

    Log entry time 00:38:26 on December 4,2008

    Entry number 251945

    keyword=SHIFT SUMMARY

    Continue production running 12uA on 3He transverse polarization 63.7%
    LHRS at 16 deg, 2.35 GeV/c, negative polarity
    Beam lamda/2 OUT
    BPMA -0.05 / 1.77
    BPMB -1.50 / 1.22

    16:00 Shift start with PRODUCTION run 4147 ongoing HAPPEX run 30499
    16:35 END RUN 4147, Q = 0.03987 C
    16:36 START RUN 4148
    17:42 END RUN 4148, Q = 0.03981 C
    17:42 START RUN 4149
    18:50 END RUN 4149, Q = 0.03979 C
    18:51 STARTED new HPPEX run 30500
    18:52 START RUN 4150
    19:07 MCC called for an RF cavity problem. Beam Down probably for more than 10 minutes.
    19:10 END RUN 4150, Q = 0.007205 ... for Beam down
    19:15 RUN 4151 during beam down taken a run for RICH pedestal
    19:24 RUN 4152 COSMICS RUN ...waiting for the beam
    20:05 an alarm the Detector Gas System Interlock Status on Low Pressure.
    Went in the gas shed and noticed that th Nitrogen bottle was empty and in alarm status.
    Also noticed that on the Hall A General Tools panel
    the Gas pressure of both Argon and CO2 report negative values
    (where ok in the day check list)
    Called Jack Segal
    20:05 MCC called. Beam is back END cosmic RUN 4152
    BPMA -0.05 / 1.74
    BPMB -1.50 / 1.20
    20:10 START Production RUN 4153
    20:30 gas back to normal condition.Jack changed the nitrogen bottles.
    Nitrogen cylinders ran out and the gas systems eventually shut down.
    21:10 observed decrease of the RICH number of photons. Pedestals analysis show many noisy channels.
    Reduced performance possibly due to a bad pedestal run
    21:15 MCC called. They have problems with two RF cavities.
    21:16 END RUN 4153 for beam down (RF cavities probloems)
    22:00 ROC 3 and 4 and then on 8,9,1 rebooted
    22:09 START RUN 4160 RICH new Pedestal during beam down (see previous comment)
    22:14 END RUN 4160
    22:25 MCC called. Problem fixed. Beam will come soon
    22:35 Beam is back at 12 uA
    BPMA 0.007 / 1.77
    BPMB -1.48 / 1.22
    22:35 Started new HAPPEX run 30501
    22:37 START 4162 production RUN
    22:43 MCC called. They take beam back to fix problems related to previous ones
    22:44 END RUN 4162 after 117k events
    23:30 MCC called ready to give us beam back
    BPM are fine
    23:34 START RUN 4163

    Accumulated production-beam charge 0.169 C